# Reflection Theory of Nichols Algebras over Coquasi-Hopf Algebras with Bijective Antipode

### v1

@@ 1.1 The classical setting: Nichols algebras over Hopf algebras
Nichols algebras play an important role in the theory of Hopf algebras and quantum groups, serving as key building blocks for the classification of pointed Hopf algebras and the structural analysis of braided tensor categories. Their combinatorial and representation-theoretic properties are intrinsically linked to the underlying Yetter–Drinfeld modules, and understanding their behavior under natural maps, such as reflections, has become a central theme in the research of Nichols algebra.
Heckenberger introduced the concepts of semi-Cartan graphs and Weyl groupoids [Heckenberger, 2006,Heckenberger, 2009,Heckenberger, 2010], which provide powerful combinatorial tools for studying the structure and classification of Nichols algebras. A semi-Cartan graph is essentially a quadruple $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},r,A)$ consisting of a finite index set $\mathbb{I}$, a set of points $\mathcal{X}$, a set of reflection maps $r$, and a generalized Cartan matrix $A^{X}$ associated with each point, satisfying specific compatibility axioms (see Definition 5.14 for details) The most celebrated theorem states that a tuple of simple Yetter-Drinfeld modules over a Hopf algebra with bijective antipode gives rise to a semi-Cartan graph, provided that it admits all reflections [Andruskiewitsch et al., 2010,Heckenberger and Schneider, 2010b,Heckenberger and Schneider, 2013a].
For Hopf algebras, the reflection theory of Nichols algebras, has been extensively developed and successfully applied to classify finite-dimensional Nichols algebras of diagonal type [Heckenberger, 2006,Heckenberger, 2010,Angiono, 2013,Angiono, 2015]. Moreover, the classification of finite-dimensional pointed Hopf algebras over abelian groups has been completed with the aid of the classification of Nichols algebras of diagonal type [Andruskiewitsch and Schneider, 1998,Andruskiewitsch and Schneider, 2000,Andruskiewitsch and Schneider, 2010,Andruskiewitsch et al., 2014,Angiono and García Iglesias, 2019,Andruskiewitsch and Angiono, 2017]. Significant progress has also been made in classifying finite-dimensional Nichols algebras over non-abelian groups [Andruskiewitsch and Graña, 1999,Andruskiewitsch and Graña, 2003,Andruskiewitsch et al., 2010,Milinski and Schneider, 2000,Fomin and Kirillov, 1999,Graña, 2000,Graña et al., 2011,Dello and Zhang, 2016,Heckenberger and Schneider, 2010b,Heckenberger and Schneider, 2013b,Heckenberger and Schneider, 2010a,Heckenberger and Vendramin, 2014,Heckenberger and Vendramin, 2015,Heckenberger and Vendramin, 2017].
@@ 1.2 From Hopf algebras to coquasi‑Hopf algebras
The primary obstacle in extending the theory to the arbitrary setting lies in the duality. We now consider two arbitrary Hopf algebras $A$, $B$ in $\mathcal{C}$ with bijective antipodes related by a non-degenerate Hopf pairing. Unfortunately, the monoidal categories $^{A}_{A}\mathcal{YD}(\mathcal{C})$} and $^{B}_{B}\mathcal{YD}(\mathcal{C})$} are not monoidal equivalence in general. More precisely, when $A$, $B$ are infinite-dimensional, a module of $A$ in $\mathcal{C}$ can not give rise to a comodule of $B$ in $\mathcal{C}$.
@@ 1.3 Braided monoidal equivalence induced by a dual pair and its application to reflection theory
In the present work, we overcome several new difficulties that do not appear in pointed cosemisimple settings and remove the assumption on finite-dimensionality. As a result, we extend the reflection theory to arbitrary coquasi‑Hopf algebras $H$ with bijective antipode.
Inspired by methods in [Heckenberger and Schneider, 2013a,Heckenberger and Schneider, [2020] ©2020], we can define rational modules of $A$ in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. This allows us to define the category ${{}^{R}_{R}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$, which is proven to be a monoidal subcategory of $^{R}_{R}\mathcal{YD}(\mathcal{C})$}. Now for two locally finite Hopf algebras $A$, $B$ in $\mathcal{C}$ with bijective antipodes related by a non-degenerate Hopf pairing compatible with the grading, called a dual pair. We prove the following result:
@@
$V_{j}=\begin{cases}M_{i}^{*},&\text{if }j=i,\\ (\mathrm{ad}\,M_{i})^{m_{ij}^{M}}(M_{j}),&\text{if }j\neq i.\end{cases}$
It is still a tuple of finite-dimensional irreducible Yetter-Drinfeld modules. By above theorem, we have a braided monoidal equivalence:
@@
(Theorem 5.12) With above assumptions on $M$. Suppose $M$ admits the $i$-th reflection for $1\leq i\leq\theta$. Then there is an isomorphism of Hopf algebras in $^{H}_{H}\mathcal{YD}$}:
@@
Finally, we apply this machinery to a crucial test case: a rank-three Nichols algebra $\mathcal{B}(M)=\mathcal{B}(M_{1}\oplus M_{2}\oplus M_{3})$ of non-diagonal type, originally identified in [Huang et al., 2026, Proposition 4.1], which is proved to be infinite-dimensional via three different approachs [Huang et al., 2026,Li and Liu, 2025a,Li and Liu, 2025b]. We compute the Weyl groupoid of $\mathcal{W}(\mathcal{G}(M))$, and show that the set of real roots of $\mathcal{G}(M)$ at each $X\in\mathcal{X}$ is just the set of real roots of Kac-Moody Lie algebra of $\mathcal{A}_{2}^{(1)}$. Furthermore, we prove that $\mathcal{G}(M)$ is a Cartan graph. Finally, utilizing the correspondence between Cartan graphs and Tits cones[Cuntz, 2018,Cuntz et al., 2019], we demonstrate that this algebra is affine, meaning its Tits cone is a half-plane. Our observations shows that affine Nichols algebras can be realized over coquasi-Hopf algebras.
@@ 1.4 Organization of this paper
The paper is organized as follows. After reviewing preliminaries on coquasi-Hopf algebras, Yetter–Drinfeld modules, and bosonization in Section 2, we introduce rational modules and establish equivalences of braided monoidal categories via dual pair in Section 3. Section 4 specializes to Nichols algebras and their pairings. Section 5 is devoted to the general reflection theory: we study the projection of a Nichols algebra and reflections of Nichols algebras. Finally, in Section 6 we work out a detailed example of an affine Nichols algebra and showing that it gives rise to a Cartan graph.
@@ 2.2 Yetter-Drinfeld module categories over coquasi-Hopf algebras
We now turn our attention to the Yetter-Drinfeld module category structure. The definition of Yetter-Drinfeld module category $^{H}_{H}\mathcal{YD}$} over an arbitrary coquasi Hopf-algebra $H$ was already given in [Ardizzoni and Pavarin, 2013]. Since it plays a crucial role in our paper, we recall this definition.
@@
The category $^{H}_{H}\mathcal{YD}$} is a $\mathbbm{k}$-linear braided monoidal abelian category over the field $\mathbbm{k}$. The unit object of $^{H}_{H}\mathcal{YD}$} is $\mathbbm{k}$, which is regarded as an object in $^{H}_{H}\mathcal{YD}$} via trivial structures. For $V,W\in{{}^{H}_{H}\mathcal{YD}}$, the tensor product of Yetter-Drinfeld modules is defined by:
@@
In Balan (2009), the authors proved that the category $\mathcal{YD}^{H}_{H}$ of right-right Yetter-Drinfeld module is braided equivalent to $\mathcal{Z}_{l}(\mathcal{M}^{H})$, which is the left center of the category of right comodules of $H$. The proof for $^{H}_{H}\mathcal{YD}$} is completely dual. We need recall the definition of right center here for completeness.
@@
Let $(V,c_{V,-})$ be an object in $\mathcal{Z}(^{H}\mathcal{M})$, we can define a left linear map on $V$ as follows:
$\rhd:H\otimes V\longrightarrow V,\qquad h\rhd v=(\mathrm{id}_{V}\otimes\varepsilon)c_{H,V}(h\otimes v)$ (2.7)
Let $W\in{{}^{H}\mathcal{M}}$ and $w^{*}\in W^{*}$. Define the map $f_{w^{*}}:W\longrightarrow H,f_{w^{*}}(w)=w_{-1}w^{*}(w_{0})$. Then $f_{w^{*}}$ is $H$-comodule map. By naturality of $c_{-,V}$, we have
$c_{H,V}\circ(f_{w^{*}}\otimes\operatorname{id}_{V})$ $=(\operatorname{id}_{V}\otimes f_{w^{*}})\circ c_{W,V}$
@@
$(\operatorname{id}_{V}\otimes w^{*})=w*(w_{0})(w_{-1}\rhd v).$
@@ 2.3 Bosonization for coquasi-Hopf algebras
$\varepsilon_{M}(r\otimes h)$ $=\varepsilon_{R}(r)\varepsilon_{M}(h),$
@@
Now suppose $N$ are $H$ are both coquasi-Hopf algebras. Furthermore, assume there exists morphisms of coquasi-Hopf algebras
@@
Given $R,R^{\prime}$ be two Hopf algebras in $^{H}_{H}\mathcal{YD}$} together with a surjective Hopf algebra map $\pi^{\prime}:R^{\prime}\rightarrow R$ and an injective map $\sigma^{\prime}:R\rightarrow R^{\prime}$ such that $\pi^{\prime}\circ\sigma^{\prime}=\operatorname{id}_{R}$. Let $M^{\prime}=R^{\prime}\#H$, and M=R#HM=R\#H. Obviously, these maps naturally extend to yield a surjective coquasi-Hopf algebra map $\pi:M^{\prime}\rightarrow M$ and an injective coquasi-Hopf algebra map $\sigma:M\rightarrow M^{\prime}$, such that $\pi\circ\sigma=\operatorname{id}_{M}$.
Let $K:=M^{\prime\operatorname{co}M}$, then $L$ is an object in $^{M}_{M}\mathcal{YD}$}, with the following structures:
$\operatorname{ad}:M\otimes K\rightarrow K,h\otimes k\mapsto\Phi_{H}(h_{1}k_{-1},\mathbb{S}_{H}(h_{3})_{1},h_{4})(h_{2}k_{0})\mathbb{S}_{H}(h_{3})_{2},$
$\operatorname{\rho}:K\rightarrow H\otimes K,k\mapsto\pi(k_{1})\otimes k_{2}.$
To analyze the linear map $\operatorname{ad}$ in concrete calculations, the following computation is useful. Let $y\in R$ be a primitive element in $^{H}_{H}\mathcal{YD}$}, One can compute the preantipode $\mathbb{S}(y)$ in $M$.
$\mathbb{S}(y)=\beta*\mathcal{S}*\alpha(y)=\beta(y_{1})\mathcal{S}_{M}(y_{2})\alpha(y_{3})=\beta(y_{-1})\mathcal{S}_{M}(y_{0})=-\beta(y_{-3})\mathcal{S}_{H}(y_{-2})\alpha(y_{-1})y_{0}.$
@@
$\operatorname{ad}(y)(x)$ $=yx+(y_{-1}x)\mathbb{S}(y_{0})$ (2.13)
$=yx+(y_{-4}x)\beta(y_{-3})(\mathcal{S}_{H}(y_{-2})\alpha(y_{-1})y_{0})$
$=yx-\Phi(y_{-7}x_{-1},\mathcal{S}(y_{-4}),y_{-1})\beta(y_{-6})((y_{-5}x_{0})\mathcal{S}(y_{-3}))\alpha(y_{-2})y_{0}.$
On the other hand, note that for all $h\in H$, $\Delta_{H}(\mathbb{S}(h))=\beta(h_{1})\mathcal{S}(h_{3})\otimes\mathcal{S}(h_{2})\alpha(h_{4})$. Therefore, we have
$yx-\operatorname{ad}(y_{-1})(x)y_{0}$ $=yx-\Phi(y_{-4}x_{-1},\mathbb{S}(y_{-2})_{1},y_{-1}))((y_{-3}x_{0})\mathbb{S}(y_{-2})_{2})y_{0}$
$=yx-\Phi(y_{-7}x_{-1},\mathcal{S}(y_{-4}),y_{-1})\beta(y_{-6})((y_{-5}x_{0})\mathcal{S}(y_{-3}))\alpha(y_{-2})y_{0}.$
@@
2.4 Yetter-Drinfeld modules in arbitrary braided monoidal category

### v2

# arXiv:2602.07946v2[math.QA] 27 Feb 2026

@@ 1.1 The classical setting: Nichols algebras over Hopf algebras
Nichols algebras play an important role in the theory of Hopf algebras and quantum groups, serving as key building blocks for the classification of pointed Hopf algebras and the structural analysis of braided tensor categories. Their combinatorial and representation-theoretic properties are intrinsically linked to the underlying Yetter–Drinfeld modules, and understanding their behavior under natural maps, such as reflections, has become a central theme in the research on Nichols algebra.
Heckenberger introduced the concepts of semi-Cartan graphs and Weyl groupoids [Heckenberger, 2006,Heckenberger, 2009,Heckenberger, 2010], which provide powerful combinatorial tools for studying the structure and classification of Nichols algebras. A semi-Cartan graph is essentially a quadruple $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},r,A)$ consisting of a finite index set $\mathbb{I}$, a set of points $\mathcal{X}$, a set of reflection maps $r$, and a generalized Cartan matrix $A^{X}$ associated with each point $X\in\mathcal{X}$, satisfying specific compatibility axioms (see Definition 5.14 for details) The most celebrated theorem states that a tuple of simple Yetter-Drinfeld modules over a Hopf algebra with bijective antipode gives rise to a semi-Cartan graph, provided that it admits all reflections [Andruskiewitsch et al., 2010,Heckenberger and Schneider, 2010b,Heckenberger and Schneider, 2013a].
For Hopf algebras, the reflection theory of Nichols algebras has been extensively developed and successfully applied to classify finite-dimensional Nichols algebras of diagonal type [Heckenberger, 2006,Heckenberger, 2010,Angiono, 2013,Angiono, 2015]. Moreover, the classification of finite-dimensional pointed Hopf algebras over abelian groups has been completed with the aid of the classification of Nichols algebras of diagonal type [Andruskiewitsch and Schneider, 1998,Andruskiewitsch and Schneider, 2000,Andruskiewitsch and Schneider, 2010,Andruskiewitsch et al., 2014,Angiono and García Iglesias, 2019,Andruskiewitsch and Angiono, 2017]. Significant progress has also been made in classifying finite-dimensional Nichols algebras over non-abelian groups [Andruskiewitsch and Graña, 1999,Andruskiewitsch and Graña, 2003,Andruskiewitsch et al., 2010,Milinski and Schneider, 2000,Fomin and Kirillov, 1999,Graña, 2000,Graña et al., 2011,Dello and Zhang, 2016,Heckenberger and Schneider, 2010b,Heckenberger and Schneider, 2013b,Heckenberger and Schneider, 2010a,Heckenberger and Vendramin, 2014,Heckenberger and Vendramin, 2015,Heckenberger and Vendramin, 2017].
@@ 1.2 From Hopf algebras to coquasi‑Hopf algebras
The primary obstacle in extending the theory to the arbitrary setting lies in the duality. We now consider two arbitrary Hopf algebras $A$, $B$ in $\mathcal{C}$ with bijective antipodes related by a non-degenerate Hopf pairing. Unfortunately, the monoidal categories $^{A}_{A}\mathcal{YD}(\mathcal{C})$} and $^{B}_{B}\mathcal{YD}(\mathcal{C})$} are not monoidal equivalent in general. More precisely, when $A$, $B$ are infinite-dimensional, a module of $A$ in $\mathcal{C}$ cannot give rise to a comodule of $B$ in $\mathcal{C}$.
@@ 1.3 Braided monoidal equivalence induced by a dual pair and its application to reflection theory
In the present work, we overcome several new difficulties that do not appear in pointed cosemisimple settings and remove the assumption of finite-dimensionality. As a result, we extend the reflection theory to arbitrary coquasi‑Hopf algebras $H$ with bijective antipode.
Inspired by methods in [Heckenberger and Schneider, 2013a,Heckenberger and Schneider, [2020] ©2020], we can define rational modules of $A$ in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. This allows us to define the category ${{}^{R}_{R}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$, which is proven to be a monoidal subcategory of $^{R}_{R}\mathcal{YD}(\mathcal{C})$}. For two locally finite Hopf algebras $A$, $B$ in $\mathcal{C}$ with bijective antipodes related by a non-degenerate Hopf pairing compatible with the grading, called a dual pair, we prove the following result:
@@
$V_{j}=\begin{cases}M_{i}^{*},&\text{if }j=i,\\ \mathrm{ad}(M_{i})^{m_{ij}^{M}}(M_{j}),&\text{if }j\neq i.\end{cases}$
It is still a tuple of finite-dimensional irreducible Yetter-Drinfeld modules. By the above theorem, we have a braided monoidal equivalence:
@@
(Theorem 5.12) Under the above assumptions on $M$. Suppose $M$ admits the $i$-th reflection for $1\leq i\leq\theta$. Then there is an isomorphism of Hopf algebras in $^{H}_{H}\mathcal{YD}$}:
@@
[…]
@@ 1.4 Organization of this paper
The paper is organized as follows. After reviewing preliminaries on coquasi-Hopf algebras, Yetter–Drinfeld modules, and bosonization in Section 2, we introduce rational modules and establish equivalences of braided monoidal categories via dual pair in Section 3. Section 4 specializes to Nichols algebras and their pairings. Section 5 is devoted to the general reflection theory: we study the projection of a Nichols algebra and reflections of Nichols algebras. Finally, in Section 6 we work out a detailed example of an affine Nichols algebra and show that it gives rise to a Cartan graph.
@@ 2.2 Yetter-Drinfeld module categories over coquasi-Hopf algebras
We now turn our attention to the Yetter-Drinfeld module category structure. The definition of the Yetter-Drinfeld module category $^{H}_{H}\mathcal{YD}$} over an arbitrary coquasi-Hopf algebra $H$ was already given in [Ardizzoni and Pavarin, 2013]. Since it plays a crucial role in our paper, we recall this definition.
@@
The category $^{H}_{H}\mathcal{YD}$} is a $\mathbbm{k}$-linear braided monoidal abelian category. The unit object of $^{H}_{H}\mathcal{YD}$} is $\mathbbm{k}$, which is regarded as an object in $^{H}_{H}\mathcal{YD}$} via trivial structures. For $V,W\in{{}^{H}_{H}\mathcal{YD}}$, the tensor product of Yetter-Drinfeld modules is defined by:
@@
In Balan (2009), the authors proved that the category $\mathcal{YD}^{H}_{H}$ of right-right Yetter-Drinfeld module is braided equivalent to $\mathcal{Z}_{l}(\mathcal{M}^{H})$, which is the left center of the category of right comodules of $H$. The proof for $^{H}_{H}\mathcal{YD}$} is completely dual. We need to recall the definition of right center here for completeness.
@@
Let $(V,c_{V,-})$ be an object in $\mathcal{Z}(^{H}\mathcal{M})$; we can define a left linear map on $V$ as follows:
$\rhd:H\otimes V\longrightarrow V,\qquad h\rhd v=(\mathrm{id}_{V}\otimes\varepsilon)c_{H,V}(h\otimes v).$ (2.7)
Let $W\in{{}^{H}\mathcal{M}}$ and $w^{*}\in W^{*}$. Define the map $f_{w^{*}}:W\longrightarrow H,f_{w^{*}}(w)=w_{-1}w^{*}(w_{0})$. Then $f_{w^{*}}$ is an $H$-comodule map. By naturality of $c_{-,V}$, we have
$c_{H,V}\circ(f_{w^{*}}\otimes\operatorname{id}_{V})(w\otimes v)$ $=(\operatorname{id}_{V}\otimes f_{w^{*}})\circ c_{W,V}(w\otimes v)$
@@
That is
$c_{H,V}(w_{-1}w^{*}(w_{0})\otimes v)=(\operatorname{id}_{V}\otimes f_{w^{*}})\circ c_{W,V}(w\otimes v).$
@@
$w^{*}(w_{0})(w_{-1}\rhd v)=(\operatorname{id}_{V}\otimes\varepsilon\circ f_{w^{*}})\circ c_{W,V}(w\otimes v)=(\operatorname{id}_{V}\otimes w^{*})\circ c_{W,V}(w\otimes v).$
@@ 2.3 Bosonization for coquasi-Hopf algebras
$\varepsilon_{M}(r\otimes h)$ $=\varepsilon_{R}(r)\varepsilon_{H}(h),$
@@
Now suppose $N$ and $H$ are both coquasi-Hopf algebras. Furthermore, assume there exist morphisms of coquasi-Hopf algebras
@@
Let $R,R^{\prime}$ be two Hopf algebras in $^{H}_{H}\mathcal{YD}$} together with a surjective Hopf algebra map $\pi^{\prime}:R^{\prime}\rightarrow R$ and an injective map $\sigma^{\prime}:R\rightarrow R^{\prime}$ such that $\pi^{\prime}\circ\sigma^{\prime}=\operatorname{id}_{R}$. Let $M^{\prime}=R^{\prime}\#H$, and M=R#HM=R\#H. Obviously, these maps naturally extend to yield a surjective coquasi-Hopf algebra map $\pi:M^{\prime}\rightarrow M$ and an injective coquasi-Hopf algebra map $\sigma:M\rightarrow M^{\prime}$, such that $\pi\circ\sigma=\operatorname{id}_{M}$.
Let $K:=M^{\prime\operatorname{co}M}$, then $K$ is an object in $^{M}_{M}\mathcal{YD}$}, with the following structures:
$\operatorname{ad}:M\otimes K\rightarrow K,h\otimes k\mapsto\Phi_{M}(h_{1}k_{-1},\mathbb{S}_{M}(h_{3})_{1},h_{4})(h_{2}k_{0})\mathbb{S}_{M}(h_{3})_{2},$
$\operatorname{\rho}:K\rightarrow M\otimes K,k\mapsto\pi(k_{1})\otimes k_{2}.$
To analyze the linear map $\operatorname{ad}$ in concrete calculations, the following computation is useful. Let $y\in R$ be a primitive element. One can compute the preantipode $\mathbb{S}_{M}(y)$ in $M$.
$\mathbb{S}_{M}(y)=\beta*\mathcal{S}_{M}*\alpha(y)=\beta(y_{1})\mathcal{S}_{M}(y_{2})\alpha(y_{3})=\beta(y_{-1})\mathcal{S}_{M}(y_{0})=-\beta(y_{-3})\mathcal{S}_{H}(y_{-2})\alpha(y_{-1})y_{0}.$
@@
$\operatorname{ad}(y)(x)$ $=yx+(y_{-1}x)\mathbb{S}_{M}(y_{0})$ (2.13)
$=yx-(y_{-4}x)\beta(y_{-3})(\mathcal{S}_{H}(y_{-2})\alpha(y_{-1})y_{0})$
$=yx-\Phi(y_{-7}x_{-1},\mathcal{S}_{H}(y_{-3}),y_{-1})\beta(y_{-6})((y_{-5}x_{0})\mathcal{S}_{H}(y_{-4}))\alpha(y_{-2})y_{0}.$
On the other hand, note that for all $h\in H$, $\Delta_{H}(\mathbb{S}(h))=\beta(h_{1})\mathcal{S}_{H}(h_{3})\otimes\mathcal{S}_{H}(h_{2})\alpha(h_{4})$. Therefore, we have
$yx-\operatorname{ad}(y_{-1})(x)y_{0}$ $=yx-\Phi(y_{-4}x_{-1},\mathbb{S}_{M}(y_{-2})_{1},y_{-1}))((y_{-3}x_{0})\mathbb{S}_{M}(y_{-2})_{2})y_{0}$
$=yx-\Phi(y_{-7}x_{-1},\mathcal{S}_{H}(y_{-3}),y_{-1})\beta(y_{-6})((y_{-5}x_{0})\mathcal{S}_{H}(y_{-4}))\alpha(y_{-2})y_{0}.$
@@
2.4 Yetter-Drinfeld modules in arbitrary braided monoidal categories